40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 29 02:33:28 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (I) optical_configuration.pngoptical_layout_ap_table2.pngoptical_layout_ap_table2.pdf
    Reply  Mon Jun 30 05:33:38 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (II) 
       Reply  Tue Jul 1 21:54:05 2008, Koji, Update, General, Re: Abs. Length Meas. setup 
    Reply  Wed Jul 2 15:14:42 2008, steve, Update, General, added beam traps 
Message ID: 590     Entry time: Sun Jun 29 02:33:28 2008     Reply to this: 599   624
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ Optical setup (I) 
I have constructed the beam injection optics for the abs length measurement.

The injection beam was coarsely aligned to the interferometer. The reflected beam from SRM was already seen at AS CCD.
I have attached the optical configration for this measurement and the optics layout at the AP table.

I am going to go to LHO for three weeks. During the absence Alberto tunes the mode matching and the alignment of the interferometer.

In the process of making this report, I noticed that one of the iris apertures is about disturbing the beam for OMCR CCD. I will check this before I go to Hanford. Also an RF spectrum analyzer is at the AP table. I try to return this near the PSL on Monday morning.

Attachment 1: Optical configuration for the abs length measurement.
1) One of the arms is locked to the PSL beam by the main control system (red).
2) A laser beam is injected from the AS port (blue). This laser essentially has different frequency from that of PSL.
3) The injected beam and the outgoing PSL beam appear at the output of the faraday in the injection system.
4) They beat each other at the frequency difference of those two lasers.
5) A PLL is used to lock the frequency difference to a local oscillator (LO).
6) The LO frequency is swept at around 3.87MHz, that is the approximate FSR frequency of the arm cavity.
7) If the LO frequency hits the FSR within the resonant width, the beating also appears at the transmitted light as the injected beam also becomes resonant to the arm cavity.
8) Amplitude of the beating at the transmitted light is measured by a RF spectrum analyzer as a function of the LO frequency. We get the FSR frequency (= the arm cavity length) from the top of the resonance.

Attachment 2: Optics at the AP table for the laser injection
700mW NPRO, laser source. vertically polarized.
Periscope, to raise the beam 1 inch to make the beam at the 4 inch elevation.
INJ_SM1/INJ_SM2, steering mirrors to align the injection beam to the IFO beam.
HWP1, half wave plate to make the beam to the farady horiz-polarized. nominal 42deg on the readout.
FI, Faraday isolator for protection of the NPRO from the returning light, for obtaining the returning light.
HWP2, to make the beam from the Faraday horiz-polarized. nominal 357deg on the readout.
MM_Lens, f=125mm to match the laser mode to the IFO beam.
SM1/SM2, steering mirrors to align the IFO beam to the Farady Isolator.
IRIS1/IRIS2, for the coarse alignment of the injection beam.
FLIP, flipper mount to turn on/off the injection optics.

Alignment procedure of the injection system
0) Ignite NPRO several hours before the experiment so that the laser frequency can be stable.
1) Turn up FLIP. Close the shutter of NPRO.
2) Adjust SM1/SM2 so that the ifo beam can appear at the output of FI.
3) Adjust height and position of IRIS1/IRIS2 with regard to the ifo beam so that the ifo beam goes through IRIS1/IRIS2 even when they are closed.
4) Turn down FLIP. Open the shutter of NPRO.
5) Adjust INJ_SM1/INJ_SM2 so that the injection beam can go through IRIS1/IRIS2 even when they are closed.
6) At this time, it is expected that the reflection of the injection beam from SRM appears at AS CCD, if SRM is aligned.
7) Adjust INJ_SM1/INJ_SM2 so that the injection beam at AS CCD can overlap to the IFO beam.
8) Confirm the beam at the output of the FI also overlaps.
---- We are here ----
9) Change the ifo configuration to the X or Y arm only.
10) Scan the crystal temperature of the 700mW NPRO in order to try to have the beating of the two beams at the PD. AS OSA may be useful to obtain the beating.
11) Once the beating is obtained, adjust INJ_SM1/INJ_SM2 such that the beating amplitude is maximized.
Attachment 1: optical_configuration.png  30 kB  | Show | Hide all | Show all
Attachment 2: optical_layout_ap_table2.png  602 kB  Uploaded Sun Jun 29 12:46:58 2008  | Hide | Hide all | Show all
optical_layout_ap_table2.png
Attachment 3: optical_layout_ap_table2.pdf  2.446 MB  Uploaded Sun Jun 29 12:49:17 2008  | Show | Hide all | Show all
ELOG V3.1.3-